OpenStudy (anonymous):

A school is using identification cards (ID cards) that consist of 3 letters selected from A to E inclusive followed by 3 digits chosen from 0 to 9 inclusive. a) How many different ID cards can be issued to students if a digit may be used more than once but all 3 letters of each ID are different? b) New ID cards are issued to all students each year and the old cards discarded. However, the old ID numbers are not used again. If, on average, the school's population increases by 10% each year and was 2000 during the year when the ID cards were first used, how many years will elapse before cards with numbers already used will have to be issued?

OpenStudy (anonymous):

I've done part A and need help with part B now.

OpenStudy (anonymous):

and I believe part B relates to creating the right equation except that I couldn't.

OpenStudy (anonymous):

Is the answer 216?? I am not sure??

OpenStudy (anonymous):

nope, it's the 15th year.

OpenStudy (anonymous):

Wait..

OpenStudy (anonymous):

in the 15th year*

OpenStudy (anonymous):

How many total numbers of ID cards can we make do you know that??

OpenStudy (anonymous):

yep, 60000.

OpenStudy (anonymous):

You are giving the answer for a part right??

OpenStudy (anonymous):

for part b I tried creating an equation of 2000+200n=60000 but it's not correct.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

I got 2000 + 200x = 43200 Ha ha ha.. Wait let check it once more..

OpenStudy (anonymous):

yeh neither urs or mine is correct.

OpenStudy (anonymous):

On right hand side we must get 5000: 2000 + 200x = 5000 200(x+10) = 5000 x = 15 But I am not getting 5000..

OpenStudy (anonymous):

why 5000? and not 60000?

OpenStudy (anonymous):

I think it is asking for numbers only.. Read the b part carefully..

OpenStudy (anonymous):

See, you have found 60000 for first part and that is right but in first part it is not the answer you will apply to all other parts.. First part is given with some condition and you have found 60000 according to that condition.. Getting??

OpenStudy (anonymous):

I got the solution...

OpenStudy (anonymous):

but shouldn't we use 60000?

OpenStudy (anonymous):

No we have to find it again according to the B part sayings..

OpenStudy (anonymous):

since that is the total number of codes that can be produced.

OpenStudy (anonymous):

Read the b part carefully dear..

OpenStudy (anonymous):

I've read it, I still don't get it.

OpenStudy (anonymous):

Read the last line: before cards with numbers already used will have to be issued? cards with numbers and not alphabet Right??

OpenStudy (anonymous):

Cards are containing three alphabets and three numbers B part is demanding for card with numbers and not card with alphabets.. Getting it or not??

OpenStudy (anonymous):

yepyep.

OpenStudy (anonymous):

Now, find the total numbers of cards now.. See, we can choose 3 numbers from 0 to 9 in 10*10*10 ways right???

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

We get 1000.. But in front there are alphabet also.. So there must be alphabets in front but we need not to arrange them here.. For arranging them you will get 60 as your answer.. But here arranging does not matter we are dealing with numbers here... So, out of 5 how many ways you can use letters in front?? I think 5 ways.. Because you can have A in first , B or C or D or E so total 5 ways.. Getting or not??

OpenStudy (anonymous):

So total cards = 5*1000 = 5000 after this you will get your answer: 2000 + 200n = 5000 get n from here..

OpenStudy (anonymous):

Able??

OpenStudy (anonymous):

I don't get ur "But here arranging does not matter we are dealing with numbers here... So, out of 5 how many ways you can use letters in front?? I think 5 ways.. Because you can have A in first , B or C or D or E so total 5 ways."

OpenStudy (anonymous):

I understand how there is 60 ways for the three spots from 5x4x3=60 and that there are 5 ways for the first slot.

OpenStudy (anonymous):

and why do you times five?

OpenStudy (anonymous):

to get 5000?

OpenStudy (anonymous):

Because the Cards have formed from 3 alphabets \(\color{green} {AND}\) 3 numbers.. And means multiply...

OpenStudy (anonymous):

I will have to go now.. I will surely explain you when I will come online.. Don't mind..

OpenStudy (anonymous):

yeh, i have to go as well, I think I will just ask my teacher the day after.

OpenStudy (anonymous):

but thanks for your help.

OpenStudy (anonymous):

Welcome.. But I couldn't lead you to your answer.. Sorry..

OpenStudy (anonymous):

it's fine.